# Causal Interventions Reveal Shared Structure Across English Filler–Gap Constructions

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as powerful sources of evidence for linguists seeking to develop theories of syntax. In this paper, we argue that causal interpretability methods, applied to LLMs, can greatly enhance the value of such evidence by helping us characterize the abstract mechanisms that LLMs learn to use. Our empirical focus is a set of English filler–gap dependency constructions (e.g., questions, relative clauses). Linguistic theories largely agree that these constructions share many properties. Using experiments based in Distributed Interchange Interventions, we show that LLMs converge on similar abstract analyses of these constructions. These analyses also reveal previously overlooked factors – relating to frequency, filler type, and surrounding context – that could motivate changes to standard linguistic theory. Overall, these results suggest that mechanistic, internal analyses of LLMs can push linguistic theory forward.
@@ 1 Introduction
Language models can generate and process utterances typically thought to require rich linguistic grammatical structure (Futrell et al., 2019; Wilcox et al., 2018; Manning et al., 2020; Hu et al., 2020), including much-studied syntactic constructions like long-distance filler–gap constructions (Wilcox et al., 2024). These results have been taken to challenge claims that these phenomena can be learned only with strong innate priors (Piantadosi, 2023; Futrell and Mahowald, 2025).
@@
Figure 1: Causal intervention overview. Here, we illustrate our methodology when we intervene within a class, transferring an embedded wh- filler–gap structure into a corresponding minimal pair that didn’t previously have one. We then show intervening across classes, inserting a wh- filler–gap into a gap-less cleft sentence
@@
Despite the strong performance, questions remain as to whether models acquire syntax in ways that are posited by linguists to be human-like (e.g., acquiring rich grammatical abstraction and syntactic structure). Causal interpretability methods now make it possible to characterize the abstract mechanisms underlying neural networks (Vig et al., 2020; Finlayson et al., 2021; Geiger et al., 2021; Meng et al., 2022; Geiger et al., 2023; Wang et al., 2023). These methods have revealed non-trivial linguistic syntactic structure is learned by models (Arora et al., 2024; Lasri et al., 2022; Finlayson et al., 2021; Mueller et al., 2022; Lakretz et al., 2019). But a key hypothesis in the history of linguistics is that seemingly different linguistic constructions can share underlying structure. For instance, compare “I wonder what the lion ate.” to “It was the gazelle that the lion ate.” The former is an embedded wh- clause and the latter is a cleft construction. These are distinct constructions but share something in common: both have a long-distance dependency with an extracted element, often specified with a linguistic trace: “I wonder whatt the lion ate ___t.” and “It was the gazellet that the lion ate ___t.” Thus, many linguistic theories predict common processing characteristics between these sentences (Fodor, 1989). On the other hand, there is also reason to expect wh- sentences to be quite different from clefts since both wh- elements and clefts have idiosyncratic properties (Ross, 1967; Culicover, 1999; Ginzburg and Sag, 2001).
@@
To tackle these questions, we take advantage of advances in large open source models as well as advances in mechanistic interpretability, specifically the Causal Abstraction framework (Geiger et al., 2023) and Distributed Alignment Search (DAS; Geiger et al. 2024). Our resulting methodology gives us direct access to the abstract causal mechanisms learned by these models. By accessing these causal mechanisms, we can take a filler-gap mechanism learned on Construction A (e.g., wh- sentences), transfer it to Construction B (e.g., clefts), and see if we get predictable filler–gap behavior (see Figure 1). If we do, this would be strong evidence of underlying shared structure learned by the model.
@@
Importantly, this method gives us a gradient measure of transfer. As such, we explore whether more similar constructions transfer more readily to each other; whether some constructions in general tend to serve as sources of transfer; whether mechanisms transfer across clauses; and whether transfer is greater when lexical items are shared across constructions (an effect predicted by the “lexical boost” in syntactic priming, whereby syntactic structures are primed more strongly when there is lexical overlap Pickering and Branigan, 1998).
@@
Ultimately, we find strong generalization in LMs across a range of filler–gap constructions, with effects observed at all positions within constructions. We observe lexical boost: effects are stronger when lexical items match (e.g., the same animacy). Moreover, we identify source constructions whose underlying mechanisms generalize broadly, as well as sink constructions that consistently benefit from such transferred mechanisms. Finally, we provide evidence that such generalization does not seem to extend across clausal boundaries.
@@
We claim these experiments make good on the promise that studying LMs can help us better understand linguistic structure and language learning in general by not just serving as proxies for data-driven learners, but by helping us develop linguistically interesting hypotheses (Futrell and Mahowald, 2025; Portelance and Jasbi, 2024; Potts, 2023). [^1] Code will be released with final version.
@@
Construction Prefix Filler NC Article NP Verb Label
Emb. Wh-Question (Know-Class) I know who/that the man liked ./him
Emb. Wh-Question (Wonder-Class) I wonder who/if the man liked ./him
Matrix Wh-Question Who/"" did the man like ?/him
Restrictive Relative Clause The boy who/and the man liked was/him
Cleft It was the boy/clear that the man liked ./the boy
Pseudo-Cleft Who/That the man liked was/it
Topicalization Actually, the boy/"" the man liked ./the boy
@@
Subject-Verb Agreement The boy/boys that the man liked is/are
Transitive/Intransitive Verbs Last night/Yesterday some/that man/boy ran/liked ./him
Table 1: Exemplar minimal pairs for each evaluated construction’s single-clause, animate extraction variant. The filler/label combinations are used to evaluate whether the model is processing the construction correctly and whether our causal interventions are successful. The final two constructions are used as control conditions. For a full set of examples, including the multi-clause and inanimate extraction variants, see Appendix A.
@@ 2 Filler–Gaps and Neural Models
Hence, filler–gaps have served as a common test-bed for LMs’ grammatical capacities. Wilcox et al. (2018) provided early positive evidence of RNN’s grammatical competence in English by comparing LMs’ surprisals for gap and gapless continuations in the presence and absence of fillers. More recently Ozaki et al. (2022) and Wilcox et al. (2024) have demonstrated LM sensitivities to linguistic constraints on these constructions. Kobzeva et al. (2023) found mixed results in Norwegian, a language known to have very different filler–gap structures and constraints than English.
@@ 3.1 Data
We design four templates per construction, differing in the extracted object’s animacy and by the number of clausal boundaries between the filler and the gap left by its extraction (one or two clauses). We manipulated animacy since changing animacy requires changing the key wh- element (“who” vs. “what”), but is not hypothesized to affect the sentence’s structure. All our templates involve the extraction of a direct object from a verb phrase and all follow a general template, allowing cross-construction alignment by position. Our general template, as well as examples of animate extraction from a single-clause variant of each construction, are in Table 1.
@@
The second control is the task of predicting a continuation after transitive or intransitive verbs. This task controls for the predicted label, ensuring that any generalization we find is meaningful, not merely due to heuristics related to the predicted labels. In order to maintain the distance between minimal contrast and prediction location, we have lexical items in faux-contrast at the filler, article, np positions, such that there is no meaningful difference in the sampled items at those positions.
@@ 3.2 Distributed Alignment Search
To localize internal mechanisms used by LMs to process our constructions of interest, we use Distributed Alignment Search (DAS; Geiger et al. 2024; Wu et al. 2023). DAS is a supervised interpretability method that can be used to assess whether a given feature is encoded in a particular set of neural activations. We rely on the 1-dimensional variant of DAS used by Arora et al. (2024). The core intervention performed is
@@
where $\mathbf{b}\in\mathbb{R}^{n}$ is a representation formed by the model when it processes a base example (right sides in Figure 1), and $\mathbf{s}\in\mathbb{R}^{n}$ is the corresponding representation formed when the model processes a source example (left sides in Figure 1). In our experiments, $\mathbf{b}$ and $\mathbf{s}$ are always the outputs of a Transformer block. Intuitively, this intervention defines a direction in the rotated feature space defined by the learned vector $\mathbf{a}\in\mathbb{R}^{n}$. This is a soft intervention targeting only the learned feature and preserving orthogonal dimensions of $\mathbf{b}$. In DAS, all LM parameters are kept frozen, and $\mathbf{a}$ is learned via a standard cross-entropy loss trained on interventions of the sort depicted in Figure 1. The goal of learning is to make the correct predictions under the intervention. For example, in the within-class intervention in Figure 1, we seek to learn an intervention that predicts a gap site (signaled by a period) even though the inputs correspond to a non-filler–gap case. The extent to which we can learn such an intervention provides the basis for assessing the hypothesis that the filler–gap dependency itself can be localized to the intervention site.
@@
We train interventions at each position from the filler onwards, and across every layer of our given LM. We use the pythia series of models (Biderman et al., 2023), a series of open-source, open-data LLMs. We run our experiments on the $1.4$, $2.8$ and $6.9$ billion parameter models. We find qualitatively similar results for all sizes, reporting those of the $1.4$b variant in the main text (results for $2.8$b and $6.9$b variants in Appendix H).
We evaluate two distinct categories of interventions: (1) single-source interventions, where for each of the $n$ constructions, $c_{i<n}$, the training dataset for DAS contains sentences sampled from the templates of $c_{i}$, and (2) leave-one-out construction interventions, where for each of the $n$ constructions, $c_{i<n}$, the training dataset contains sentences sampled from the templates of $c_{j\not=i}$ – that is, all constructions that are not ‘left-out’.
@@
For evaluation, we use the odds metric from Arora et al. (2024). This metric measures how much more likely a counterfactual label is after performing an intervention, with higher odds denoting larger causal effect from the given intervention. Intuitively, it tells us: after intervention, how much more likely is the continuation expected based on the “source sentence” than the one naively expected based on the “base sentence”. We measure the average odds at each position-layer pair across 400 sentences, sampled from the templates of each individual construction.
In cases of aggregation, we max-pool the average odds value across layers at each sentential position (we refer to this metric as max odds hereafter). We also normalize the max odds by the corresponding average max odds for the items present in the training set, with this normalization giving us a measure of how much the mechanisms used by a given set of constructions generalize to an evaluated construction, relative to how much they generalize to those they were trained on. We aggregate across layers by max-pooling odds because our causal methodology aims to localize syntactic features in the model, with the maximum value representing the most causally efficacious localization of the given features.
Figure 2: Average normalized max odds across positions, ±$\pm 1$ standard error. Corresponding multi-clause plots can be found in Appendix E. Note that normalization fixes the “Same Animacy, In Train Set” condition at 1.00.
Figure 3: For each source construction, we measure the odds at each position layer pair, aggregating the values by evaluation group. Corresponding plots with control values and multi-clause variants are in Appendix E.
@@
We measure the max odds for all trained interventions evaluated on every construction of the same clausal category (for a discussion on cross-clause generalization see Section 6). We then group these values into six categories, depending on the relation between the set of constructions the interventions were trained on and those used to generate the evaluation set. These groups comprise (1) the same set of constructions in the training set and the evaluation set, with the same animacy – this is our reference group as training and evaluation sentences are drawn from the same distribution; (2) the same set of constructions in training set and evaluation set, with different animacy; (3) evaluation on the held-out constructions, but with the same animacy as the training set; (4) evaluation on the held-out constructions, and differing animacy from the training set; and (5–6) the two controls.
@@
We hypothesize that the max odds for all our targeted evaluation groups will be greater than that of the controls. We further expect max odds to be higher when the evaluated constructions are in training or match in animacy.
@@ Results
Figure 2 shows the average max odds of the aforementioned groups at each position in our single-clause templates. In both these single-clause variants and the embedded-clause variants of our constructions (corresponding figure in Appendix E), we find consistently high max odds values for each of the aforementioned non-control groups. The controls show significantly less transfer. We run pairwise t-tests with a Holm-Bonferroni correction, finding the max odds of each of our test groups is significantly higher than both controls at every position in the single clause templates and nearly every position in the multi-clause ones. These results strongly suggest shared internal representations across filler–gap constructions in the evaluated models.
To test our hypotheses regarding the effect of training and evaluation set overlap and matching animacy, we fit a linear mixed-effects regression model to our max odds data at each position. Our random effects comprise intervention training set, and evaluation construction, and our fixed effects take the form of binary indicator variables for (1) whether the evaluated construction was in the training set and (2) whether the animacy condition of the evaluated construction matches that of the training set. We find significant, positive, effects for overlap, matching animacy, and their interaction at the filler, the, and np positions, and for matching animacy at the verb position. See Section C.1 for regression details. Thus, across positions, LM internal processing is sensitive to linguistically meaningful features, such as animacy of the extracted item (possible evidence of “lexical boost”).
While we broadly see generalization as fitting into held-out constructions (Figure 3), embedded wh- questions and restrictive relative clauses show noticeably less generalization than other constructions. We briefly offer up two accounts for these peculiarities: (1) there is asymmetry in LM generalization between different filler–gap dependencies or (2) these constructions are processed by largely different mechanisms than the other constructions. Clarifying which of these applies to each construction helps motivate our next experiment.
@@
Our previous experiment demonstrated significant overlap between the LM’s abstract representations of various filler–gap constructions. However, we also observed notable variation in the strength of this generalization across positions and constructions. Here, we attempt to characterize the nature of this cross-construction generalization. In particular, we attempt to identify whether there exist constructions which serve as sources (their filler–gap properties transfer well to other constructions) or sinks (filler–gap properties from other constructions transfer well to them). We further investigate which features of natural language (e.g. distributional properties construction frequency, or linguistic properties like the nature of the filler item) may drive this generalization.
@@
To characterize the degree to which a given construction is a source or sink, we perform the following procedure. First, we evaluate all single-source interventions on all constructions of the same clausal length, averaging the normalized max odds across the animacy-conditions at each position, training construction, and evaluation construction triple. We take the resulting $n\times n$ matrix to be an adjacency matrix for a weighted, directed graph $G=(V,E)$ in which vertex $V$ is a construction and each directed edge $E_{i,j}$ is the transfer from construction $i$ to construction $j$. We then calculate the out-degree centrality – the fraction of a graph’s total a given node’s outgoing edges are connected to – and in-degree centrality – the fraction of nodes its incoming edges come from. We do this for nodes (constructions) across a range of edge thresholds – that is, the minimum edge weight retained in the graph. We measure each construction’s area under the threshold-centrality curves (AUC). The resulting out- and in-degree AUCs serve as proxies for the degree to which a given construction is a source or sink respectively. We provide a representative generalization network (for the the position) in Figure 4. That figure shows particularly strong transfer into pseudoclefts, very little transfer into either control, strong within-construction tranfser (dark recurrent arrows), and some non-random structure of transfer across constructions.
Figure 4: Top: Generalization network at single-clause the position with edge-threshold of 1. Nodes size proportional to in-degree, edge size and color proportional to odds of the source construction’s interventions measured on the target construction. Bottom: In- and out-degree centrality AUCs against construction frequency.
We also analyzed the effect of construction frequency on generalization capacity. We extracted estimates of each construction’s prevalence in the English-EWT Universal Dependencies dataset (De Marneffe et al., 2021; Nivre et al., 2020; Silveira et al., 2014). See Appendix D for details.
We further investigate the effects of four parameters of linguistic variation across filler–gap constructions: the nature of the filler, whether the head daughter is inverted, the syntactic category of the mother (the word under which a construction is embedded), and the semantic/pragmatic nature of the construction (whether the fronted element is fronted by necessity or for discourse reasons). At each position, we fit a linear mixed-effects model with binary indicator variables denoting whether the source and evaluated construction match for each of the above posited parameters of variation as fixed effects, with random effects for training-source construction, and evaluated construction. For regression details, see Section C.2.
@@
We expect to see specific constructions serving as strong sources and others as strong sinks in the generalization network. We further expect a positive relationship between a given construction’s frequency and the degree to which it is a source, and conversely, a negative relationship between its frequency and its sink-ness. Finally, we anticipate stronger generalization between linguistically similar constructions than dissimilar ones.
@@
Figure 4 shows construction frequency against in-degree and out-degree AUCs, mean-pooled across sentence positions. Constructions are spread across the AUC axis, suggesting varying levels of generalization. These AUCs are consistent across both sentence position and clausal variant (single and multi-clause AUCs, faceted by position are available in Appendix F).

### v2

@@ Abstract
Language Models (LMs) have emerged as powerful sources of evidence for linguists seeking to develop theories of syntax. In this paper, we argue that causal interpretability methods, applied to LMs, can greatly enhance the value of such evidence by helping us characterize the abstract mechanisms that LMs learn to use. Our empirical focus is a set of English filler–gap dependency constructions (e.g., questions, relative clauses). Linguistic theories largely agree that these constructions share many properties. Using experiments based in Distributed Interchange Interventions, we show that LMs converge on similar abstract analyses of these constructions. These analyses also reveal previously overlooked factors – relating to frequency, filler type, and surrounding context – that could motivate changes to standard linguistic theory. Overall, these results suggest that mechanistic, internal analyses of LMs can push linguistic theory forward.
[Uncaptioned image]   https://github.com/SashaBoguraev/causal-filler-gap
@@ 1 Introduction
Language models can generate and process utterances typically thought to require rich linguistic grammatical structure (Futrell et al., 2019; Wilcox et al., 2018; Manning et al., 2020; Hu et al., 2020), including much-studied syntactic constructions like long-distance filler–gap constructions (Wilcox et al., 2024). These results have been taken to challenge claims that these phenomena can be learned only with strong innate priors (Piantadosi, 2024; Futrell and Mahowald, 2025).
@@
Figure 1: Causal intervention overview. Here, we illustrate our methodology when we intervene within a class, transferring an embedded wh- filler–gap structure into a corresponding minimal pair that didn’t previously have one. We then show intervening across classes, inserting a wh- filler–gap into a gap-less cleft sentence.
@@
Despite the strong performance, questions remain as to whether models acquire syntax in ways that are posited by linguists to be human-like (e.g., acquiring rich grammatical abstraction and syntactic structure). Causal interpretability methods now make it possible to characterize the abstract mechanisms underlying neural networks (Vig et al., 2020; Finlayson et al., 2021; Geiger et al., 2021; Meng et al., 2022; Geiger et al., 2023; Wang et al., 2023). These methods have revealed non-trivial linguistic syntactic structure is learned by models (Lakretz et al., 2019; Finlayson et al., 2021; Mueller et al., 2022; Lasri et al., 2022; Arora et al., 2024). But a key hypothesis in the history of linguistics is that seemingly different linguistic constructions can share underlying structure. For instance, compare “I wonder what the lion ate.” to “It was the gazelle that the lion ate.” The former is an embedded wh- clause and the latter is a cleft construction. These are distinct constructions but share something in common: both have a long-distance dependency with an extracted element, often specified with a linguistic trace: “I wonder whatt the lion ate ___t.” and “It was the gazellet that the lion ate ___t.” Thus, many linguistic theories predict common processing characteristics between these sentences (Fodor, 1989). On the other hand, there is also reason to expect wh- sentences to be quite different from clefts since both wh- elements and clefts have idiosyncratic properties (Ross, 1967; Culicover, 1999; Ginzburg and Sag, 2001).
@@
To tackle these questions, we take advantage of advances in large open source models as well as in mechanistic interpretability, specifically the Causal Abstraction framework (Geiger et al., 2023) and Distributed Alignment Search (DAS; Geiger et al. 2024). Our resulting methodology gives us direct access to the abstract causal mechanisms learned by these models. By accessing these causal mechanisms, we can take a filler-gap mechanism learned on Construction A (e.g., wh- sentences), transfer it to Construction B (e.g., clefts), and see if we get predictable filler–gap behavior (see Figure˜1). If we do, this would be strong evidence of underlying shared structure learned by the model.
@@
Importantly, this method gives us a gradient measure of transfer. As such, we explore whether more similar constructions transfer more readily to each other; whether some constructions in general tend to serve as sources of transfer; whether mechanisms transfer across clauses; and whether transfer is greater when lexical items are shared across constructions (an effect predicted by the “lexical boost” in syntactic priming, whereby syntactic structures are primed more strongly when there is lexical overlap; Pickering and Branigan, 1998).
@@
Ultimately, we find strong generalization in LMs across a range of filler–gap constructions, with effects observed at all positions within constructions. We observe lexical boost: effects are stronger when lexical items match (e.g., the same animacy). We find greater transfer between constructions that share linguistically relevant features (e.g., the nature of the filler or whether fronting alters the utterance’s information structure). Moreover, we identify source constructions whose underlying mechanisms generalize broadly, as well as sink constructions that consistently benefit from such transferred mechanisms. Finally, we provide evidence that such generalization does not seem to extend across clausal boundaries.
@@
We claim these experiments make good on the promise that studying LMs can help us better understand linguistic structure and language learning in general by not just serving as proxies for data-driven learners, but by helping us develop linguistically interesting hypotheses (Potts, 2023; Futrell and Mahowald, 2025).
@@
Construction Prefix Filler NC Article NP Verb Label Filler Inverted Embed Front
Emb. Wh-Q (Know) I know who/that the man liked ./him Wh No VP No
Emb. Wh- (Wonder) I wonder who/if the man liked ./him Wh No VP No
Matrix Wh-Q Who/"" did the man like ?/him Wh Yes N/A No
Restr. Rel. Clause The boy who/and the man liked was/him Wh No NP No
Cleft It was the boy/clear that the man liked ./the boy Null No VP Yes
Pseudo-Cleft Who/That the man liked was/it Wh No N/A Yes
Topicalization Actually, the boy/"" the man liked ./the boy Phrase No N/A Yes
@@
Subject-Verb Agr. The boy/boys that the man liked is/are minimal non-filler-gap control
Trans/Intrans Once/Today some/that man/boy ran/liked ./him lexically matched control
Table 1: Left Block: Exemplar minimal pairs for each construction’s single-clause, animate extraction variant (and controls). The filler/label combinations are used to evaluate whether the model is processing the construction correctly and whether our causal interventions are successful. NC (‘no comparison’) shows extra words required for the grammaticality of some constructions. As many don’t require them, we do not train or test on them. For full sets of examples, including multi-clause and inanimate extraction variants, see Appendix˜A. Right Block: Parameters of linguistic variation for the same constructions. Our parameters (columns) are (1) nature of the filler (the class of the item which has ‘filled’ the gap); (2) syntactic-head child inversion (whether the child of the constructions’ syntactic-head has inverted to appear linearly before it); (3) syntactic category of the parent (the phrase under which the construction is embedded), and (4) the semantic/pragmatic nature of the construction (whether element fronting is done by syntactic necessity or for discourse purposes). Related regression results are in Table˜2.
@@ 2 Filler–Gaps and Neural Models
Hence, filler–gaps have served as a common test-bed for LMs’ grammatical capacities. Wilcox et al. (2018) provided early positive evidence of RNNs’ grammatical competence in English by comparing LMs’ surprisals for gap and gapless continuations in the presence and absence of fillers. More recently Ozaki et al. (2022) and Wilcox et al. (2024) have demonstrated LM sensitivities to linguistic constraints on these constructions. Kobzeva et al. (2023) found mixed results in Norwegian, a language known to have very different filler–gap structures and constraints than English.
@@ 3.1 Data
We design four templates per construction, differing in the extracted object’s animacy and by the number of clausal boundaries between the filler and the gap left by its extraction (one or two clauses). We manipulated animacy since changing animacy requires changing the key wh- element (“who” vs. “what”), but is not hypothesized to affect the sentence’s structure. All our templates involve the extraction of a direct object from a verb phrase and all follow a general template, allowing cross-construction alignment by position. Our general template, as well as examples of animate extraction from a single-clause variant of each construction, can be found in Table 1.
@@
The second control is the task of predicting a continuation after transitive or intransitive verbs. This task controls for the predicted label, ensuring that any generalization we find is meaningful, not merely due to heuristics related to the predicted labels. In order to maintain the distance between minimal contrast and prediction location, we have lexical items in faux-contrast at the filler, article, and np positions, such that there is no meaningful difference in the sampled items at those positions.
@@ 3.2 Distributed Alignment Search
To localize internal mechanisms used by LMs to process our constructions of interest, we use Distributed Alignment Search (DAS; Wu et al. 2023; Geiger et al. 2024). DAS is a supervised interpretability method that can be used to assess whether a given feature is encoded in a particular set of neural activations. We rely on the 1-dimensional variant of DAS used by Arora et al. (2024). The core intervention performed is
@@
where $\mathbf{b}\in\mathbb{R}^{n}$ is a representation formed by the model when it processes a base example (right sides in Figure˜1), and $\mathbf{s}\in\mathbb{R}^{n}$ is the corresponding representation formed when the model processes a source example (left sides in Figure˜1). In our experiments, $\mathbf{b}$ and $\mathbf{s}$ are always the outputs of a Transformer block. Intuitively, this intervention defines a direction in the rotated feature space defined by the learned vector $\mathbf{a}\in\mathbb{R}^{n}$. This is a soft intervention targeting only the learned feature and preserving orthogonal dimensions of $\mathbf{b}$. In DAS, all LM parameters are kept frozen, and $\mathbf{a}$ is learned via a standard cross-entropy loss trained on interventions of the sort depicted in Figure˜1. The goal of learning is to make the correct predictions under the intervention. For example, in the within-class intervention in Figure˜1, we seek to learn an intervention that predicts a gap site (signaled by a period) even though the inputs correspond to a non-filler–gap case. The extent to which we can learn such an intervention provides the basis for assessing the hypothesis that the filler–gap dependency itself can be localized to the intervention site.
@@
We train interventions at each position from the filler onwards, and across every layer of our given LM. We use the pythia series of models (Biderman et al., 2023), a series of open-source, open-data LMs. We run our experiments on the $1.4$, $2.8$ and $6.9$ billion parameter models. We find qualitatively similar results for all sizes, reporting those of the 1.41.4b variant in the main text (results for 2.82.8b and 6.96.9b variants in Appendix˜H).
We train two distinct categories of interventions: (1) single-source interventions, where for each of the $n$ constructions, $c_{i<n}$, the training dataset for DAS contains sentences sampled from the templates of $c_{i}$, and (2) leave-one-out interventions, where for each of the $n$ constructions, $c_{i<n}$, the training dataset contains sentences sampled from the templates of $c_{j\not=i}$ – that is, all constructions that are not ‘left-out’. [^1] See Rodriguez et al. (2025) for a similar transfer approach to study semantic property inheritance in models. In both settings, our training sets consist of 200 sentences sampled from the relevant constructions, before adding each sentence’s minimal pair, resulting in perfectly balanced training sets of 400 sentences.
@@
For evaluation, we use the odds metric from Arora et al. (2024). This metric measures how much more likely a counterfactual label (i.e., the mismatched word) is after performing an intervention, with higher odds denoting larger causal effect from the given intervention. Intuitively, it tells us: after intervention, how much more likely is the continuation expected based on the “source sentence” than the one naively expected based on the “base sentence”. For each construction, we measure the average odds at each position-layer pair across 400 sentences, sampled so as to ensure no overlap with our training sets.
In cases of aggregation, we max-pool the average odds value across layers at each position (we refer to this metric as max odds hereafter). We also normalize the max odds by the corresponding average max odds for the items present in the training set. This normalization measures how much the mechanisms used by a given set of constructions generalize to an evaluated construction, relative to how much they generalize to those they were trained on. We aggregate across layers by max-pooling odds because our methodology aims to localize syntactic features in the model, with the maximum value representing the most causally efficacious localization of the given features.
Figure 2: Average normalized max odds across positions, $\pm 1$ standard error. Corresponding multi-clause plots can be found in Appendix E. Note that normalization fixes the “Same Animacy, In Train Set” condition at 1.00.
Figure 3: For each source construction, we measure the odds at each position–layer pair, aggregating the values by evaluation group. Corresponding plots with control values and multi-clause variants are in Appendix E.
@@
We measure the max odds for all trained interventions evaluated on every construction of the same clausal category (for a discussion on cross-clause generalization see Section˜6). We then group these values into six categories, based on the relation between the set of constructions the interventions were trained on and those used to generate the evaluation set. These groups comprise (1) the same set of constructions in the training set and the evaluation set, with the same animacy – this is our reference group as training and evaluation sentences are drawn from the same distribution; (2) the same set of constructions in training set and evaluation set, with different animacy; (3) evaluation on the held-out constructions, but with the same animacy as the training set; (4) evaluation on the held-out constructions, and differing animacy from the training set; and (5–6) the two controls. Conditions (1) and (2) are examples of the ‘single source interventions’ as described in Section˜3.2, with the single construction present in the training set also being present in the evaluation set, with (3) and (4) examples of ‘leave-one-out interventions’, with training sets including all constructions except the one present in the evaluation set. The sole difference between the items within these pairs is whether the animacy conditions match in the training and evaluation sets.
@@
We hypothesize that the max odds for all our targeted evaluation groups will be greater than that of the controls. We further expect max odds to be higher when the evaluated constructions are in the training set or match in animacy.
@@ Results
Figure˜2 shows the average max odds of the aforementioned groups at each position in our single-clause templates. In both these single-clause variants and the multi-clause variants of our constructions (corresponding figure in Appendix˜E), we find consistently high max odds values for each of the non-control groups. The controls show significantly less transfer. We run pairwise t-tests with a Holm-Bonferroni correction, finding the max odds of each of our test groups is significantly higher than both controls at every position in the single-clause templates and nearly every position in the multi-clause ones. These results strongly suggest shared internal representations across filler–gap constructions in the evaluated models.
To test our hypotheses regarding the effect of training and evaluation set overlap and matching animacy, we fit a linear mixed-effects regression model to our max odds data at each position. Our random effects are intervention training set and evaluation construction, and our fixed effects take the form of binary indicator variables for (1) whether the evaluated construction was in the training set and (2) whether the animacy condition of the evaluated construction matches that of the training set. We find significant, positive, effects for overlap, matching animacy, and their interaction at the filler, the, and np positions, and for matching animacy at the verb position. See Section˜C.1 for regression details. Thus, across positions, LM internal processing is sensitive to linguistically meaningful features, such as animacy of the extracted item (possible evidence of “lexical boost”).
While we broadly see generalization as fitting into held-out constructions (Figure˜3), embedded wh- questions and restrictive relative clauses show noticeably less generalization than other constructions. We briefly offer up two accounts for these peculiarities: (1) there is asymmetry in LM generalization between different filler–gap dependencies or (2) these constructions are processed by largely different mechanisms than the other constructions. Clarifying which of these applies to each construction helps motivate our next experiment.
@@
Our previous experiment demonstrated significant overlap between the LM’s abstract representations of various filler–gap constructions. However, we also observed notable variation in the strength of this generalization across positions and constructions. Here, we aim to characterize the nature of this cross-construction generalization. In particular, we aim to identify whether there exist constructions which serve as sources (their filler–gap properties transfer well to other constructions) or sinks (filler–gap properties from other constructions transfer well to them). We further investigate which features of natural language (e.g. distributional properties like frequency, or linguistic properties like the nature of the filler item) drive this generalization.
@@
To characterize the degree to which a given construction is a source or sink, we perform the following procedure. First, we evaluate all single-source interventions on all constructions of the same clausal length, averaging the normalized max odds across the animacy-conditions at each position, training construction, and evaluation construction triple. At each position, we take the resulting $n\times n$ matrix to be an adjacency matrix for a weighted, directed graph $G=(V,E)$ in which vertex $V$ is a construction and each directed edge $E_{i,j}$ is the transfer from construction $i$ to construction $j$. We then calculate out-degree centrality – the fraction of a graph’s total nodes that a given node’s outgoing edges are connected to – and in-degree centrality – the fraction of nodes its incoming edges come from. We do this across a range of edge thresholds – i.e., the minimum edge weight retained in the graph. We measure each construction’s area under the threshold-centrality curves (AUC). The resulting out- and in-degree AUCs serve as proxies for the degree to which a given construction is a source or sink respectively. We provide an exemplar generalization network (for the the position) in Figure˜4. That figure shows particularly strong transfer into pseudo-clefts, very little transfer into either control, strong within-construction transfer (dark recurrent arrows), and some non-random structure of transfer across constructions.
Figure 4: Top: Generalization network at single-clause the position with edge-threshold of 1. Node size proportional to in-degree; edge size and color proportional to odds of the source construction’s interventions measured on the target construction. Bottom: In- and out-degree centrality AUCs against construction frequency.
We also analyzed the effect of construction frequency on generalization capacity. We extracted estimates of each construction’s prevalence in the English-EWT Universal Dependencies dataset (De Marneffe et al., 2021; Nivre et al., 2020; Silveira et al., 2014). See Appendix˜D for details.
We further investigate the effects of four parameters of linguistic variation across filler–gap constructions: the nature of the filler, whether the head child is inverted, the syntactic category of the parent (the word under which a construction is embedded), and the semantic/pragmatic nature of the construction (whether the fronted element is fronted by necessity or for discourse reasons), with Table˜1 presenting the associated parameter value for each construction. At each position, we fit a linear mixed-effects model predicting the max odds, with binary indicator variables denoting whether the source and evaluated construction match for each of the above posited parameters of variation as fixed effects, with random effects for training-source construction, and evaluated construction. For full regression details, see Section˜C.2.
Finally, we perform Principal Component Analysis (PCA) at each position, reducing the dimensionality of our generalization matrix to the two principal components, allowing visualization of construction similarity in this space.
@@
We expect some constructions to serve as strong sources and others as strong sinks in the generalization network. We further expect a positive relationship between a construction’s frequency and the degree to which it is a source, and conversely, a negative relationship between its frequency and its sink-ness. Finally, we anticipate stronger generalization between linguistically similar constructions than dissimilar ones (as operationalized by the parameters of linguistic variation in Table˜1).
@@
Figure 5: Constructions plotted along the top two principal components at each position in our single-clause variants. Generally, constructions cluster in linguistically intuitive ways – e.g. animate/inanimate pairs generally cluster, constructions with wh-fillers cluster at the filler position, and restrictive relative clauses typically lie away from the other analyzed constructions.
Figure˜4 shows construction frequency against in-degree and out-degree AUCs, mean-pooled across sentence positions. Constructions are spread across the AUC axis, suggesting varying levels of generalization. These AUCs are consistent across both sentence position and clausal variant (single and multi-clause AUCs, faceted by position, are available in Appendix˜F).